OpenStudy (anonymous):

what single discount is equivalent to two succesive discounts of 15% and 7%?

OpenStudy (anonymous):

20.95

OpenStudy (anonymous):

I do believe that if you just combine the two percentages they are equal to two discounts back to back.

OpenStudy (anonymous):

100-100*15/100=85 85-85*7/100=79.05 100-79.05=20.95

OpenStudy (anonymous):

thank you for your help!